brusslesprout:

i need help with this one question on my math homework. "Due to recent demand, a store raised the price of a video game by 24% the game previously sold for $25. What is the new price?"

Rx1nyDxyz:

Ill help

Rx1nyDxyz:

What do you think you have to do?

brusslesprout:

I have an example of a previous problem if that might help? 63/84 = x/100 x = 75

Rx1nyDxyz:

Yeah, you're heading in the right direction.

Rx1nyDxyz:

Ok, when you're increasing, you multiply the original price by what?

Rx1nyDxyz:

Are you there?

brusslesprout:

I'm here. My apologies

Rx1nyDxyz:

That's alright. Anyway, what do you think you multiply the original price by?

Rx1nyDxyz:

If you don't know, i'll help you out. It's by the 1+ percent of increase. For this problem, it would be 1.24

Rx1nyDxyz:

So, what's 25 x 1.24

Rx1nyDxyz:

That will give you your answer.

brusslesprout:

31?

Rx1nyDxyz:

Exactly! You got this :)

brusslesprout:

It's stated as one of the possible answers on my homework. Thank you!

Rx1nyDxyz:

Yw :) Tag me if you need any help